# MnLargeSymbols’164 MnLargeSymbols’171

### v2

@@ 4 Solution to the Mean Field Game
$C(\mu,\nu,\mu_{0},T)=24T^{2}\left(\frac{\Theta}{\theta^{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}^{-1}\right)\max\left\{P_{2}(\mu)^{\frac{1}{2}},P_{2}(\nu)^{\frac{1}{2}}\right\}+\frac{\Theta^{2}}{\theta^{2}}\right)^{2}\mathbb{E}\left[h_{0}^{2}\right]e^{TC_{2,2}(T,\nu(\cdot))+6TL_{f}^{2}\frac{\Theta^{2}}{\theta^{2}}M(\mu)^{2}+3T\zeta^{2}+3\chi^{2}}.$ (48)
@@
where $|\cdot|_{[0,T]\times(-r,r)\times(0,r)}^{(\beta)}$ denotes the Hölder seminorm of exponent $\beta$ in $[0,T]\times(-r,r)\times(0,r)$, $\beta>0$, and $C(T,r,\mu)$ depends on $T,r,\overline{M}(\mu(\cdot))$ and is independent of $n$.
@@
$-\partial_{t}W_{n}+\rho W_{n}=\tilde{H}_{1}(t,x,y,\mu_{n}(t),D_{y}W_{n})+\frac{1}{2}\chi^{2}D^{2}_{yy}W_{n}+H_{0}(D_{x}W_{n})$ (53)
@@
[…]
@@
Estimates (49) and (50) can also be proved directly, with computations very similar to those in the proofs of Propositions 3.2 and 3.3, thanks to the fact that for every fixed sequence $\mu_{n}(\cdot)$ one can bound $\overline{M}(\mu_{n}(\cdot))$ with a constant independent of $n$.
@@
First we prove that $V_{n}(t_{0},\bar{x}_{0},\bar{h}_{0})\rightarrow V(t_{0},\bar{x}_{0},\bar{h}_{0})$ for all $t_{0}\in[0,T],\bar{x}_{0}\in\mathbb{R},\bar{h}_{0}\in\mathbb{R}_{+}$; we will write $\bar{\delta}_{0}$ for $\delta_{\bar{x}_{0},\bar{h}_{0}}$ (the Dirac measure with mass $1$ in $(\bar{x}_{0},\bar{h}_{0})$ and mass 00 everywhere else). For all $n\in\mathbb{N}$ and $\varepsilon>0$, let $(v_{n},s_{n})\in\mathcal{K}$ be $\varepsilon$-optimal controls for $V_{n}(t_{0},x_{0},h_{0})$, denote by $\left(x_{n}(\cdot),h_{n}(\cdot)\right)$ the solution to
@@
$\left|V_{n}(t_{0},\bar{x}_{0},\bar{h}_{0})-V_{\mu}(t_{0},\bar{x}_{0},\bar{h}_{0})\right|\leq\varepsilon+\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|J_{n}(x_{n}(t),h_{n}(t),s_{n}(t),v_{n}(t))-J_{\mu}(x_{n}(t),h_{\mu}(t),s_{n}(t),v_{n}(t))\right|\mathrm{d}t\right],$
@@
[…]
@@
and the same conclusion actually holds for every measure $\nu\in\mathcal{P}_{2}$ in place of $\bar{\delta}_{0}$). By the arbitrariness of $\varepsilon$ we deduce the convergence of $V_{n}$ to $V$, locally uniformly on $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$.
By Lemma 4.2, $D_{h}V_{n}$ is uniformly bounded in $n$ and uniformly Hölder continuous in $n$ in every compact set of $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$. Set $K_{1}=[0,T]\times(-r,)]\times(0,r)$ for ome fixed $r>0$. By the Ascoli-Arzelà theorem we can extract a subsequence $\left\{D_{h}V_{n}^{1}\right\}$ of $\left\{D_{n}V_{n}\right\}$ which converges locally uniformly on $K_{1}$. Similarly, we set $K_{2}=[0,T]\times(-r-1,r+1)\times(0,r+1)$ and we can extract a subsequence $\left\{D_{h}V_{n}^{2}\right\}$ of $\left\{D_{h}V_{n}^{1}\right\}$ that converges locally uniformly $K_{2}$. We repeat the same argument on each relatively compact set $K_{k}=[0,T]\times(-r-k+1,r+k-1)\times(0,r+k-1)$ and deduce the existence, for each $k$, of a subsequence $\left\{D_{h}V_{n}^{k}\right\}$ uniformly converging in $K_{k}$. Then, the diagonal subsequence $\left\{D_{h}V_{m}^{m}\right\}$ converges locally uniformly to some function $Q$. But also $V_{n}$ converges to $V$ locally uniformly, so that by the regularity of $V_{n}$ and $V$ we must have $Q=D_{h}V$. Since this construction can be applied to every subsequence of $\{D_{h}V_{n}\}$, also the original sequence $\{D_{h}V_{n}\}$ converges locally uniformly to $D_{h}V$. The exact same argument can be applied to $\{D_{x}V_{n}\}$, yielding its locally uniform convergence to $D_{x}V$. By the arbitrariness of $(t_{0},\bar{x}_{0},\bar{h}_{0})$ we deduce the convergence of $V_{n}$, $D_{x}V_{n}$ and $D_{h}V_{n}$ on $[0,T]\times\mathbb{R}\times\mathbb{R}_{++}$.
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(t,x,h)m_{n_{k}}(t;\mathrm{d}x,\mathrm{d}h)\rightarrow\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(t,x,h)\bar{m}(t;\mathrm{d}x,\mathrm{d}h).$
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{0}$ $(D_{x}V_{n_{k}}(r,x,h))D_{x}\phi(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{0}(D_{x}V(r,x,h))D_{h}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
@@
+×$\phantom{\leq}+\int_{\mathbb{R}\times\mathbb{R}_{+}}\left(D_{p}H_{0}(D_{x}V_{n_{k}}(r,x,h))-D_{p}H_{0}(D_{x}V(r,x,h))\right)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
@@
for some constant $\hat{L}$, because $D_{x}V_{n}$ is bounded uniformly in $n$ and $D_{p}H_{0}$ is locally Lipschitz. The right hand side in the previous inequality then converges to 00 because $m_{n_{k}}$ converges weakly to $\bar{m}$ and $D_{x}V_{n_{k}}$ converges locally uniformly to $D_{x}V$.
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{\partial\phi}{\partial t}(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{\partial\phi}{\partial t}(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}D^{2}_{xx}\phi(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}D^{2}_{xx}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
trivially converge to 00 since $m_{n_{k}}$ converges to $\bar{m}$ weakly.
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}h^{2}D^{2}_{hh}\phi(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}h^{2}D^{2}_{hh}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
converges to 00.
@@
×$-\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{1}(x,h,\bar{m}(r),D_{h}V(r,x,h))D_{h}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$ (54)
×$\phantom{-}\leq\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))D_{h}\phi(r,x,h)\left(m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\bar{m}(r;\mathrm{d}x,\mathrm{d}h)\right)$ (55)
+×$\phantom{-\leq}+\int_{\mathbb{R}\times\mathbb{R}_{+}}\left(D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))\right.$ (56)
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))D_{h}\phi(r,x,h)\left(m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\bar{m}(r;\mathrm{d}x,\mathrm{d}h)\right)\\ \leq\hat{C}_{\phi}\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{h}\phi(r,x,h)\left(m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\bar{m}(r;\mathrm{d}x,\mathrm{d}h)\right);$
therefore (55) converges to 00. By Lemma 3.5 we have that $D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))-D_{p}H_{1}(x,h,\bar{m}(r),D_{h}V(r,x,h))$ converges to 00 pointwise, while being uniformly bounded for the same reason as above; therefore by the dominated convergence theorem also (56) converges to 00.

### v3

@@ 4 Solution to the Mean Field Game
$C(\mu,\nu,\mu_{0},T)=24T^{2}\left(\frac{\Theta}{\theta^{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}^{-1}\right)\max\left\{P_{2}(\mu)^{\frac{1}{2}},P_{2}(\nu)^{\frac{1}{2}}\right\}+\frac{\Theta^{2}}{\theta^{2}}\right)^{2}\\ \cdot\mathbb{E}\left[h_{0}^{2}\right]e^{TC_{2,2}(T,\nu(\cdot))+6TL_{f}^{2}\frac{\Theta^{2}}{\theta^{2}}M(\mu)^{2}+3T\zeta^{2}+3\chi^{2}}.$ (48)
@@
where $|\cdot|_{[0,T]\times(-r,r)\times(0,r)}^{(\beta)}$ denotes the Hölder seminorm of exponent $\beta$ in $0,T]\times(-r,r)\times(0,r)$, $\beta>0$, and $C(T,r,\mu)$ depends on $T,r,\overline{M}(\mu(\cdot))$ and is independent of $n$.
@@
$\partial_{t}W_{n}+\rho W_{n}=\tilde{H}_{1}(t,x,y,\mu_{n}(t),D_{y}W_{n})+\frac{1}{2}\chi^{2}D^{2}_{yy}W_{n}+H_{0}(D_{x}W_{n})$ (53)
@@
$|W_{n}(t_{0},x_{0},y_{0})|=|V_{n}(t_{0},x_{0},e^{y_{0}})|$ $\leq$ $C\overline{M}(\mu_{n}(\cdot))^{\gamma(1-\sigma)}\int_{t_{0}}^{T}e^{-\rho t}\mathbb{E}\left[h_{n}(t)^{\eta}\right]\,dt$
$\leq$ $C\overline{M}(\mu_{n}(\cdot))^{\gamma(1-\sigma)}\int_{t_{0}}^{T}e^{-\rho t}\mathbb{E}\left[h_{n}(t)\right]^{\eta}\,dt$
@@
Estimates (49) and (50) can also be proved directly, with computations very similar to those in the proofs of Propositions 3.3 and 3.4, thanks to the fact that for every fixed sequence $\mu_{n}(\cdot)$ one can bound $\overline{M}(\mu_{n}(\cdot))$ with a constant independent of $n$.
@@
First we prove that $V_{n}(t_{0},\bar{x}_{0},\bar{h}_{0})\rightarrow V(t_{0},\bar{x}_{0},\bar{h}_{0})$ for all $t_{0}\in[0,T],\bar{x}_{0}\in\mathbb{R},\bar{h}_{0}\in\mathbb{R}_{+}$; we will write $\bar{\delta}_{0}$ for $\delta_{\bar{x}_{0},\bar{h}_{0}}$ (the Dirac measure with mass $1$ in $(\bar{x}_{0},\bar{h}_{0})$ and mass 0 everywhere else). For all $n\in\mathbb{N}$ and $\varepsilon>0$, let $(v_{n},s_{n})\in\mathcal{K}$ be $\varepsilon$-optimal controls for $V_{n}(t_{0},x_{0},h_{0})$, denote by $\left(x_{n}(\cdot),h_{n}(\cdot)\right)$ the solution to
@@
$\left|V_{n}(t_{0},\bar{x}_{0},\bar{h}_{0})-V_{\mu}(t_{0},\bar{x}_{0},\bar{h}_{0})\right|\leq\\ \varepsilon+\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|J_{n}(x_{n}(t),h_{n}(t),s_{n}(t),v_{n}(t))-J_{\mu}(x_{n}(t),h_{\mu}(t),s_{n}(t),v_{n}(t))\right|\mathrm{d}t\right],$
@@
$\leq\varepsilon+\hat{C}d_{\infty,2}(\mu_{n}(\cdot),\mu(\cdot))^{\gamma}\int_{t_{0}}^{T}\mathbb{E}[h_{n}(t)^{2}]^{\frac{\eta}{2}}dt+\hat{C}\int_{t_{0}}^{T}\mathbb{E}[\left|h_{n}(t)-h_{\mu}(t)\right|^{2}]^{\frac{\eta}{2}}dt$
@@
and the same conclusion actually holds for every measure $\nu\in\mathcal{P}_{2}$ in place of $\bar{\delta}_{0}$). By the arbitrariness of $\varepsilon$ we deduce the convergence of $V_{n}$ to $V$, locally uniformly on $0,T]\times\mathbb{R}\times\mathbb{R}_{+}$.
By Lemma 4.2, $D_{h}V_{n}$ is uniformly bounded in $n$ and uniformly Hölder continuous in $n$ in every compact set of $0,T]\times\mathbb{R}\times\mathbb{R}_{+}$. Set $K_{1}=[0,T]\times(-r,)]\times(0,r)$ for ome fixed $r>0$. By the Ascoli-Arzelà theorem we can extract a subsequence $\left\{D_{h}V_{n}^{1}\right\}$ of $\left\{D_{n}V_{n}\right\}$ which converges locally uniformly on $K_{1}$. Similarly, we set $K_{2}=[0,T]\times(-r-1,r+1)\times(0,r+1)$ and we can extract a subsequence $\left\{D_{h}V_{n}^{2}\right\}$ of $\left\{D_{h}V_{n}^{1}\right\}$ that converges locally uniformly $K_{2}$. We repeat the same argument on each relatively compact set $K_{k}=[0,T]\times(-r-k+1,r+k-1)\times(0,r+k-1)$ and deduce the existence, for each $k$, of a subsequence $\left\{D_{h}V_{n}^{k}\right\}$ uniformly converging in $K_{k}$. Then, the diagonal subsequence $\left\{D_{h}V_{m}^{m}\right\}$ converges locally uniformly to some function $Q$. But also $V_{n}$ converges to $V$ locally uniformly, so that by the regularity of $V_{n}$ and $V$ we must have $Q=D_{h}V$. Since this construction can be applied to every subsequence of $\{D_{h}V_{n}\}$, also the original sequence $\{D_{h}V_{n}\}$ converges locally uniformly to $D_{h}V$. The exact same argument can be applied to $\{D_{x}V_{n}\}$, yielding its locally uniform convergence to $D_{x}V$. By the arbitrariness of $(t_{0},\bar{x}_{0},\bar{h}_{0})$ we deduce the convergence of $V_{n}$, $D_{x}V_{n}$ and $D_{h}V_{n}$ on $0,T]\times\mathbb{R}\times\mathbb{R}_{++}$.
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(t,x,h)m_{n_{k}}(t;\mathrm{d}x,\mathrm{d}h)\rightarrow\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(t,x,h)\bar{m}(t;\mathrm{d}x,\mathrm{d}h).$
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{0}$ $(D_{x}V_{n_{k}}(r,x,h))D_{x}\phi(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)$
$\phantom{\leq}-\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{0}(D_{x}V(r,x,h))D_{h}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
@@
$\phantom{\leq}+\int_{\mathbb{R}\times\mathbb{R}_{+}}\left(D_{p}H_{0}(D_{x}V_{n_{k}}(r,x,h))-D_{p}H_{0}(D_{x}V(r,x,h))\right)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
@@
for some constant $\hat{L}$, because $D_{x}V_{n}$ is bounded uniformly in $n$ and $D_{p}H_{0}$ is locally Lipschitz. The right hand side in the previous inequality then converges to 0 because $m_{n_{k}}$ converges weakly to $\bar{m}$ and $D_{x}V_{n_{k}}$ converges locally uniformly to $D_{x}V$.
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{\partial\phi}{\partial t}(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{\partial\phi}{\partial t}(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}D^{2}_{xx}\phi(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}D^{2}_{xx}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
trivially converge to 0 since $m_{n_{k}}$ converges to $\bar{m}$ weakly.
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}h^{2}D^{2}_{hh}\phi(r,x,h)m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}h^{2}D^{2}_{hh}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$
converges to 0.
@@
$\phantom{aaaa}-\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{1}(x,h,\bar{m}(r),D_{h}V(r,x,h))D_{h}\phi(r,x,h)\bar{m}(r;\mathrm{d}x,\mathrm{d}h)$ (54)
$\phantom{-}\leq\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))D_{h}\phi(r,x,h)\left(m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\bar{m}(r;\mathrm{d}x,\mathrm{d}h)\right)$ (55)
$\phantom{-\leq}+\int_{\mathbb{R}\times\mathbb{R}_{+}}\left(D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))\right$. (56)
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))D_{h}\phi(r,x,h)\left(m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\bar{m}(r;\mathrm{d}x,\mathrm{d}h)\right)\\ \leq\hat{C}_{\phi}\int_{\mathbb{R}\times\mathbb{R}_{+}}D_{h}\phi(r,x,h)\left(m_{n_{k}}(r;\mathrm{d}x,\mathrm{d}h)-\bar{m}(r;\mathrm{d}x,\mathrm{d}h)\right);$
therefore (55) converges to 0. By Lemma 3.6 we have that $D_{p}H_{1}(x,h,m_{n_{k}}(r),D_{h}V_{n_{k}}(r,x,h))-D_{p}H_{1}(x,h,\bar{m}(r),D_{h}V(r,x,h))$ converges to 0 pointwise, while being uniformly bounded for the same reason as above; therefore by the dominated convergence theorem also (56) converges to 0.